# Who is better at math, Jenny or Jingzhen? Uncovering Stereotypes in Large Language Models

### v2

@@ 1 Introduction
Existing research on stereotypes in LMs is limited, and predominantly focuses on African American and White groups Jiang and Fellbaum (2020); May et al. (2019), or a subset of US census groups, often with Middle Eastern added Guo and Caliskan (2020); Cao et al. (2022); Kirk et al. (2021); Cheng et al. (2023). Furthermore, datasets that seek to expand the coverage of bias measures to multiple axes are limited to a fixed set of stereotypes for specific demographic groups Nangia et al. (2020); Nadeem et al. (2021); Parrish et al. (2022). To address these limitations, we focus on incorporating a wide range of ethnicities, allowing us to gain a more international view of the effects of bias. We also highlight the importance of analysis that uses a intersectional lens, where biases compound across a combination of different axes, e.g., gender and ethnicity, to cause unique harms.
@@
    the GlobalBias [^1] The dataset and code used to evaluate the models can be found at https://github.com/groovychoons/GlobalBias. dataset for studying harmful stereotypes, which consists of 876,000 sentences for 40 distinct gender-by-ethnicity groups
@@ 2.3 Stereotype Datasets
This paper builds on previous work using stereotypes as a means of exploring characteristics that are observed to be associated with specific demographic groups, reinforcing established social hierarchies Greenwald et al. (1998); Blodgett et al. (2021). A number of datasets have been developed to look specifically at stereotypes, largely focusing on using sentence pairs to compare two demographic groups May et al. (2019), a stereotype and anti-stereotype Zhao et al. (2018a); Nangia et al. (2020); Nadeem et al. (2021), or a question and answer set to compare two groups Parrish et al. (2022). While valuable, these datasets are all situated within a U.S. context Blodgett et al. (2021), and are unsuitable for our context of analysing multiple subgroups across a single stereotype. In contrast, our one-vs-all approach offers more robust statistical power, and reduces the impact of outliers and natural variability in two-group comparisons. In terms of metrics, we posit that perplexity, which Smith et al. (2022) and Smith and Williams (2021) used to compare multiple subgroups in one test, is a suitable method, and thus we develop it further.
@@ 3 The Dataset
We propose a new dataset named GlobalBias for studying harmful stereotypes, which consists of sets of 10 proper names spanning 40 groups. A summary of key data can be found in Table 2.
@@ 3.1 Proper Names
Our seed dataset of names is the Genni + Ethnea dataset Torvik (2018). It contains over 2 million names, each annotated with ethnicity and gender. We first filter $\sim$176,000 unique first names, to include only those with 2 to 14 characters and a male or female gender classification, narrowing our dataset to approximately 35,000 names. We exclusively included names labeled with a binary gender by the Genni model used to label the seed dataset, thereby excluding gender-neutral names. We posit that gender-neutral names do not necessarily represent gender diverse groups in LLMs, and are more often a mixture of male and female stereotypes, though we acknowledge that focusing on binary gender classification fails to represent the diverse spectrum of human self-identification as discussed in Butler (1989) and Kuper et al. (2012).
By utilizing embeddings and clustering techniques, we identify names that an LLM perceives as highly correlated within these groups. We use OpenAI’s (text-embedding-3-large) embeddings for each name and apply Mini Batch K-Means clustering to group the names into clusters. We select ten names per group to prevent one name from having a large impact on results. These ten names are randomly selected from clusters with a high gender and ethnicity agreement, i.e. $>50\%$ names in the same group in a cluster, meaning an LLM is likely to classify the chosen cluster as belonging to that ethnic and gender group. Where ethnicities have an exclusive gender, we select 10 names of the opposite gender with high probability of belonging to that ethnicity for gender balance across the entire dataset.
@@ 3.2 Descriptors
We initially draw on three existing datasets: the HOLISTICBIAS dataset Smith et al. (2022), Ghavami and Peplau (2013) and StereoSet Nadeem et al. (2021). The first, HOLISTICBIAS, is split into 13 demographic axes; we use 11 of these axes (Race/Ethnicity and Nationality are excluded, as the purpose of the experiment is to infer these from the given name). Ghavami and Peplau (2013) provides a labelled dataset of stereotypes from a free-response survey. We extract stereotype terms from StereoSet, which was handcrafted to test a fixed set of stereotypes in LLMs. As a result, our descriptor terms represent a diverse range of potential stereotypes.
@@ 4.3 Models
In our experiments, we evaluate a suite of seven language models to examine the generalizability of our bias measures across various model sizes and architectures, these are: BERT (google-bert/bert-large-cased; Devlin et al. 2018), RoBERTa (roberta-large; Liu et al. 2019), Flan-UL2 (google/flan-ul2, Tay et al. 2023), GPT-2 (gpt2-xl, Radford et al. 2019), GPT Neo X (EleutherAI/gpt-neox-20b; Black et al. 2022), OPT (facebook/opt-30b; Zhang et al. 2022) and Llama 3 (meta-llama/Meta-Llama-3-8B; AI@Meta 2024).
@@ 6 Stereotypes via Generation
6.1 Classification
To assess the the level of bias in each model, we construct a one-vs-all SVM classification across gender, ethnicity, and gender-by-ethnicity groups, to measure how easily differentiable demographic groups are from each other. We partition our data in to 70% for training and 30% for testing, stratified based on demographic group. Each character profile was represented using 11 features, with each feature encoded as either a one-hot vector (for single words) or sparse vector of the relative frequencies of the words in the feature (for lists of words).
@@ 6.2 Feature Analysis
We conduct a feature elimination process to identify the importance of different features in distinguishing between demographic groups, in order to identify potential sources of bias. We analyse groups of features such as ’hobbies’, rather than individual features such as ’reading’. A full table of the impact of each group of features can be found in Appendix C.
@@ 7 Conclusion
We find that larger models produce more stereotypical outputs, even when explicitly instructed not to. We also show using the example of religion that bias stays consistent across model’s internal representation and outputs. Through this investigation, we aim to raise awareness about the importance of considering intersectionality when evaluating models and encourage nuanced and thoughtful evaluation of stereotypes in LLMs.
@@
While our work aims to broaden the scope of ethnicities covered in NLP bias research, there are many ethnic groups and genders not covered in this work, and we exclude other critical aspects such as age, disability, and socioeconomic status. The dataset’s creation process excludes gender-neutral names, limiting its applicability to a broader spectrum of identities. Moreover, the GlobalBias dataset is not intended as a benchmark; instead, it is used to gain insights into a wider set of intersectional demographic groups.
@@ Appendix B Generation Prompt
“hobbies” - list of strings, length 3
@@
“socioeconomic_status” - string
“sexual_orientation” - string
“religion” - string
@@ Appendix C Feature Analysis in Generation
Feature Eliminated Llama 3 70B GPT 3.5 Claude 3 Opus GPT 4o
Overall Accuracy (%) 18.3% 21.7% 26.4% 33.3%
age +2.8% $\uparrow$ +0.5% $\uparrow$ +0.8% $\uparrow$ +0.6% $\uparrow$
@@ Appendix D Top Words in Generation
sexual_orientation lesbian gay asexual bisexual Slav Female, Dutch Male, Turkish Male, Israeli Female French Male, African Female Japanese Female Japanese Male, Italian Female

### v3

@@ 1 Introduction
Existing research on stereotypes in LMs is limited, and predominantly focuses on African American and White groups Jiang and Fellbaum (2020); May et al. (2019), or a subset of US census groups, often with Middle Eastern added Guo and Caliskan (2021); Cao et al. (2022); Kirk et al. (2024); Cheng et al. (2023). Furthermore, datasets that seek to expand the coverage of bias measures to multiple axes are limited to a fixed set of stereotypes for specific demographic groups Nangia et al. (2020); Nadeem et al. (2021); Parrish et al. (2022). To address these limitations, we focus on incorporating a wide range of ethnicities and using a one-vs-all, unsupervised approach to identify which stereotypes are associated with each demographic group. We also highlight the importance of analysis that uses a intersectional lens, where biases compound across a combination of different axes, e.g., gender and ethnicity, to cause unique harms.
@@
    the GlobalBias dataset for studying harmful stereotypes, which consists of 876,000 sentences for 40 distinct gender-by-ethnicity groups
@@ 2.3 Stereotype Datasets
This paper builds on previous work exploring how stereotypes are associated with specific demographic groups and how this reinforces existing social hierarchies Greenwald et al. (1998); Blodgett et al. (2021). Several datasets have been developed to examine stereotypes, often structured around sentence pairs comparing two demographic groups May et al. (2019), contrasting stereotypes and anti-stereotypes Zhao et al. (2018a); Nangia et al. (2020); Nadeem et al. (2021), or using question-answer sets to compare groups Parrish et al. (2022). While valuable, these datasets are not suitable for our objective of analyzing each stereotype across multiple subgroups. In contrast, our one-vs-all approach offers more robust statistical power, and reduces the impact of outliers and natural variability in two-group comparisons. For evaluation, we posit that perplexity, which Smith et al. (2022) and Smith and Williams (2021) used to compare multiple subgroups in one test, is a suitable method, and thus we develop it further.
Furthermore, the datasets mentioned above are predominantly situated within a U.S. context Blodgett et al. (2021), and do not adequately represent the global user base of LLMs. While efforts have been made to adapt these datasets for other languages and cultural contexts Fort et al. (2024); Sahoo et al. (2024), our methodology bypasses the need for pre-defined stereotype labels, allowing for more flexible analysis of outputs across various contexts.
@@ 3 The Dataset
We propose a new dataset [^1] The dataset and code used to evaluate the models can be found at https://github.com/groovychoons/GlobalBias. named GlobalBias for studying harmful stereotypes, which consists of sets of 10 proper names spanning 40 groups. A summary of key statistics for the dataset can be found in Table 2.
@@ 3.1 Proper Names
Our seed dataset of names is the Genni + Ethnea dataset Torvik (2018). It contains over 2 million names, each annotated with ethnicity and gender. We first filter $\sim$176,000 unique first names, to include only those with 2 to 14 characters and a male or female gender classification, narrowing our dataset to approximately 35,000 names. We exclusively included names labeled with a binary gender by the Genni model used to label the seed dataset, thereby excluding gender-neutral names. We posit that gender-neutral names do not necessarily represent gender diverse groups in LLMs, and are more often a mixture of male and female stereotypes, though we acknowledge that focusing on binary gender classification fails to represent the diverse spectrum of human self-identification as discussed in Butler (1989), Kuper et al. (2012) and Larson (2017).
By utilizing embeddings and clustering techniques, we identify names that an LLM perceives as highly correlated within these groups. We use OpenAI’s (text-embedding-3-large) embeddings for each name and apply Mini Batch K-Means clustering to group the names into clusters. We select ten names per group to prevent one name from having a large impact on results, and reduce the harm caused by misclassification of names Gautam et al. (2024). These ten names are randomly selected from clusters with a high gender and ethnicity agreement, i.e. $>50\%$ names in the same group in a cluster, meaning an LLM is likely to classify the chosen cluster as belonging to that ethnic and gender group. Where ethnicities have an exclusive gender, we select 10 names of the opposite gender with high probability of belonging to that ethnicity for gender balance across the entire dataset.
@@ 3.2 Descriptors
We initially draw on three existing datasets: the HOLISTICBIAS dataset Smith et al. (2022), Ghavami and Peplau (2013) and StereoSet Nadeem et al. (2021). The first, HOLISTICBIAS, is split into 13 demographic axes; we use 11 of these axes (Race/Ethnicity and Nationality are excluded, as the purpose of the experiment is to infer these from the given name). Ghavami and Peplau (2013) provide a labelled dataset of stereotypes from a free-response survey. We extract stereotype terms from StereoSet, which was handcrafted to test a fixed set of stereotypes in LLMs. As a result, our descriptor terms represent a diverse range of potential stereotypes.
@@ 4.3 Models
In our experiments in Sections 4 and 5, we evaluate a suite of seven language models to examine the generalizability of our bias measures across various model sizes and architectures, these are: BERT (google-bert/bert-large-cased; Devlin et al. 2019), RoBERTa (roberta-large; Liu et al. 2019), Flan-UL2 (google/flan-ul2, Tay et al. 2022), GPT-2 (gpt2-xl, Radford et al. 2019), GPT Neo X (EleutherAI/gpt-neox-20b; Black et al. 2022), OPT (facebook/opt-30b; Zhang et al. 2022) and Llama 3 (meta-llama/Meta-Llama-3-8B; AI@Meta 2024).
@@ 6 Stereotypes via Generation
6.1 Classification
To assess the level of bias in each model, we construct a one-vs-all SVM classification across gender, ethnicity, and gender-by-ethnicity groups, to measure how easily differentiable demographic groups are from each other. We partition our data in to 70% for training and 30% for testing, stratified based on demographic group. Each character profile was represented using 11 features, with each feature encoded as either a one-hot vector (for single words) or sparse vector of the relative frequencies of the words in the feature (for lists of words).
Feature Eliminated Llama 3 70B GPT 3.5 Claude 3 Opus GPT 4o
Overall Accuracy (%) 18.3% 21.7% 26.4% 33.3%
age +2.8% $\uparrow$ +0.5% $\uparrow$ +0.8% $\uparrow$ +0.6% $\uparrow$
@@ 6.2 Feature Analysis
We conduct a feature elimination process to identify the importance of different features in distinguishing between demographic groups, in order to identify potential sources of bias. We analyse groups of features such as ‘hobbies’, rather than individual features such as ‘reading’. The impact of each group of features for gender-by-ethnicity groups can be found in Table 7. The impact of each group of features for ethnicity only and gender only can be found in Appendix C.
@@ 7 Conclusion
We find that larger models produce more stereotypical outputs, even when explicitly instructed not to. We also show using the example of religion that bias stays consistent across model’s internal representation and outputs.
Our work reveals the prevalence and impact of stereotypes across a diverse range of ethnic and gender groups through the introduction of the GlobalBias dataset. We highlight the importance of a comprehensive and intersectional approach to studying bias in LMs, which is essential for ensuring ethical, fair, and effective use of LMs in real-world scenarios, ultimately fostering trust and inclusivity in technology.
@@
While our work aims to broaden the scope of ethnicities covered in NLP bias research, there are many ethnic groups and genders not covered in this work, and we exclude other critical aspects such as age, disability, and socioeconomic status. The dataset’s creation process excludes gender-neutral names, limiting its applicability to a broader spectrum of identities, and that the use of given names in itself can contribute to harm Gautam et al. (2024). We encourage future data collection involving given names to allow self-identification of gender, where possible, as recommended by Larson (2017). Moreover, the GlobalBias dataset is not intended as a benchmark; instead, it is used to gain insights into a wider set of intersectional demographic groups.
@@ Appendix B Generation Prompt
"hobbies" - list of strings, length 3
@@
"socioeconomic_status" - string
"sexual_orientation" - string
"religion" - string
@@ Appendix D Top Words in Generation
sexual_orientation lesbian gay asexual bisexual Slav Female, Dutch Male, Turkish Male, Israeli Fem. French Male, African Female Japanese Female Japanese Male, Italian Female